# 1Manifold Research 2Metarch.ai 3Georgia Tech

### v1

@@ IV-C1 Patterns of Collapse Influenced by Decoding Techniques
Diffusion models inherently differ from AR models due to their generative approach, which aims to reconstruct data from noise by modeling the entire distribution of training data [34]. In multi-class classification tasks with mutually exclusive single-correct answers, diffusion models respond distinctly when encountering OOD inputs. Primarily, diffusion models excel at reconstructing data similar to their training distribution. However, upon encountering OOD inputs, their reconstruction accuracy deteriorates markedly due to difficulty in mapping the unfamiliar inputs onto the learned data manifold. Consequently, predictions on OOD data are characterized by uncertainty, with predictions distributed across multiple classes instead of sharply peaked at a single class. Visual inspection of confusion matrices from Pi0 Base on datasets like Bigfish, Climber, and Maze as seen in Figures 24, 25, and 26, supports this finding, demonstrating distributed prediction frequencies rather than dominance by any one class.
@@
In contrast, AR models sequentially predict outputs conditioned on previous outputs, typically applying softmax at the final step for classification tasks. A known limitation of this approach is an inherent tendency toward overconfidence in OOD scenarios. Despite inputs being significantly divergent from training data, AR models frequently produce sharply peaked probability distributions, arbitrarily favoring one class. This behavior arises from their inductive biases, where the softmax activation overestimates probabilities for certain classes. This issue is evidenced clearly in the union confusion matrices seen in Figure 11. Each AR model consistently collapses its predictions onto particular classes, even with nonsensical or significantly divergent inputs, as observed in the individual per-subdataset confusion matrices in the Appendix section VII-A for OpenVLA’s outcomes on Chaser and Jumper, Pi0 FAST’s results on Bigfish, Chaser, and Heist, and GPT 4x’s performance on Heist, Maze, Coinrun and Miner.
@@
The prediction collapse behavior also varies significantly due to differences in tokenization methods employed by AR models. Specifically, OpenVLA and Pi0 FAST differ fundamentally in their tokenization and decoding strategies, influencing their OOD performance distinctly.
@@
Pi0 FAST utilizes the Discrete Cosine Transform (DCT) [1] in combination with Byte Pair Encoding (BPE) [64], focusing on low-frequency components that represent overarching shape of the signals. This frequency-domain approach, particularly the strategy of flattening DCT coefficients prioritizing low-frequency information, results in concentrated predictions due to stronger global pattern capture and robust representation against input noise. This approach enhances consistency and stability in action predictions, as evidenced by concentrated class predictions observed in the confusion matrices.
@@
Conversely, OpenVLA employs a discrete binning approach, mapping continuous action spaces to discrete bins directly, thus not capturing global patterns in any sense when evaluated in a zero-shot setting. Each action dimension or timestep is processed independently, which leads to a more diffused spread of predictions. OpenVLA exhibits sensitivity to variations and input noise, resulting in more evenly distributed probability mass across action classes, as reflected in the broader spread observed in its confusion matrices.
@@ V-1 Differences Between Training Data and Procgen
Dataset Domain Discrepancies
A fundamental obstacle limiting model performance arises from substantial differences between the training datasets and the Procgen evaluation dataset. Procgen environments are procedurally generated, Atari-like 2D games designed to evaluate visual and motor skills of reinforcement learning (RL) agents. Each subdataset within Procgen exhibits diverse environment layouts, tasks, objectives, reward structures, and discrete action spaces, predominantly involving directional movements and specific game-related interactions.
@@
OpenVLA is a vision-language model specifically fine-tuned on robotics data from the Open X-Embodiment dataset, comprising approximately 970,000 robot demonstrations. These demonstrations are strictly limited to continuous robotic manipulation tasks recorded from third-person camera views.
Pi0 models also leverage robotics trajectory data featuring continuous action spaces. Their training incorporates a subset of OpenX-Embodiment alongside a private dataset containing approximately 903 million timesteps, significantly sourced from single and dual-arm robotic manipulation scenarios.
@@
Action Space Awareness
Each model exhibits varying degrees of awareness regarding the action space constraints:
GPT 4x models are explicitly informed through prompts about valid action values and corresponding verbal descriptions.
@@
Pi0 Base and Pi0 Fast rely on training-induced normalization but lack explicit clamping, occasionally resulting in invalid predictions outside permissible action ranges.
Proprioceptive State Handling
Procgen datasets lack proprioceptive state information, a critical input for Pi0 models during training. To compensate, zero arrays replace missing proprioceptive states during inference, negatively impacting Pi0 model performance by introducing unnatural, non-representative states. While OpenVLA optionally accepts proprioceptive states, the absence of these in Procgen data inherently limits the richness of information available for inference.
Image Input Variability
Significant discrepancies in image inputs further exacerbate model adaptation difficulties:
@@
View differences: OpenVLA expects a single image input, matching the Procgen input structure, potentially enhancing its adaptability. In contrast, Pi0 models trained on multiple views suffer due to missing additional viewpoints, where zero arrays substitute these inputs, further degrading performance.
@@ V-2 Difficulty Adapting to Increased Image Complexity
We quantified image complexity through Shannon entropy [66] and Delentropy [57] across the first 20 episodes per dataset as seen in Figures 13 and 14, revealing distinctive performance patterns. The formula used to calculate the entropies can be seen in Appendix section VII-B:
Figure 13: Entropy measures VS Model Macro Recall
Figure 14: Correlation matrix for Entropy measures VS Model Macro Recall
GPT-4 models (4o and 4.1) demonstrated moderate to weak negative correlations (Shannon entropy: -0.41 and -0.28; Delentropy: -0.351 and -0.25, respectively), suggesting poorer performance with increasing complexity.
OpenVLA and Pi0 Base exhibited negligible correlations with entropy metrics (OpenVLA: 0.071 Shannon, 0.17 Delentropy; Pi0 Base: 0.047 Shannon, 0.063 Delentropy), implying relative independence from image complexity.
Pi0 Fast uniquely displayed moderate positive correlations (Shannon and Delentropy: 0.44), indicating improved performance with higher image complexity.
@@ V-3 Impact of VLA Training, Decoding, and Output Processing Techniques
Pi0 Base also targets normalized outputs within [-1,1] without enforced clamping. Its diffusion-based architecture inherently predicts actions ”close” to familiar training distributions. This approach yields relatively fewer invalid outputs and moderately effective generalization compared to purely autoregressive models but remains inferior to OpenVLA.
@@ VI Conclusion
In this study, we systematically evaluated the generalization capabilities of contemporary VLAs and VLMs across procedurally generated discrete-action environments from the Procgen dataset in a zero-shot setting. Our analysis highlighted significant limitations arising from architectural constraints, training paradigms, and input-output biases inherent to the models. The stark domain discrepancy between training data—primarily continuous-action robotics datasets, and general web-scale vision language data—and discrete-action game environments emerged as a critical barrier to effective zero-shot generalization.
We identified notable differences in model behaviors linked directly to their architectures, training strategies, and input/output processing techniques. OpenVLA’s robust action-space clamping technique consistently provided superior generalization, minimizing invalid outputs and exhibiting relative resilience to out-of-distribution scenarios. Conversely, autoregressive models like GPT-4x displayed substantial difficulties in generalizing, especially under complex image conditions, and frequently defaulted to overly simplistic or biased action choices. Additionally, Pi0 models showed intermediate performance influenced heavily by their diffusion-based (Pi0 Base) or autoregressive (Pi0 FAST) decoding methods, with Pi0 FAST notably sensitive to image complexity and unable to restrict majority of its predictions to a desired output range.
Our findings underscore the necessity for architectural innovations, refined training methodologies, and enhanced output processing techniques to bridge the gap between diverse action domains. Future research should prioritize developing more generalized training datasets that better reflect the variety of potential application environments, alongside methods to adaptively handle different forms of action representations. These advancements hold promise for enabling VLAs to operate effectively across an increasingly diverse and unpredictable range of real-world tasks.

### v2

@@ IV-C1 Patterns of Collapse Influenced by Decoding Techniques
Diffusion models inherently differ from AR models due to their generative approach, which aims to reconstruct data from noise by modeling the entire distribution of training data [15]. In multi-class classification tasks with mutually exclusive single-correct answers, diffusion models respond distinctly when encountering OOD inputs. Primarily, diffusion models excel at reconstructing data similar to their training distribution. However, upon encountering OOD inputs, their reconstruction accuracy deteriorates markedly due to difficulty in mapping the unfamiliar inputs onto the learned data manifold. Consequently, predictions on OOD data are characterized by uncertainty, with predictions distributed across multiple classes instead of sharply peaking at a single class. Visual inspection of the confusion matrices from Pi0 Base on datasets like Bigfish, Climber, and Maze as seen in Figures 34, 35, and 36, supports this finding, demonstrating distributed prediction frequencies rather than dominance by any one class.
@@
In contrast, AR models sequentially predict outputs conditioned on previous outputs, typically applying softmax at the final step for classification tasks. A known limitation of this approach is an inherent tendency toward overconfidence in OOD scenarios. Despite inputs being significantly divergent from training data, AR models frequently produce sharply peaked probability distributions, arbitrarily favoring one class. This behavior arises from their inductive biases, where the softmax activation overestimates probabilities for certain classes. This issue is evidenced clearly in the union confusion matrices seen in Figure 19 and the action distributions seen in 14, 15, 16, 17, and 18. Each AR model consistently collapses its predictions onto particular classes, even with nonsensical or significantly divergent inputs, as observed in the individual per-subdataset confusion matrices in the Appendix section VII-B for OpenVLA’s outcomes on Chaser and Jumper, Pi0 FAST’s results on Bigfish, Chaser, and Heist, and GPT 4x’s performance on Heist, Maze, Coinrun and Miner.
The difference in collapse behaviors between the AR and Diffusion models can be further observed by comparing the action distributions of Pi0 Base and Pi0 FAST in Figure 21, as both models share the PaliGemma VLM backbone in their architecture.
Figure 21: Pi0 Base vs. Pi0 FAST action distributions.
@@
The prediction collapse behavior also varies significantly between different AR models due to differences in tokenization methods. Specifically, OpenVLA and Pi0 FAST differ fundamentally in their tokenization and decoding strategies, influencing their OOD performance distinctly.
@@
Pi0 FAST utilizes the Discrete Cosine Transform (DCT) [1] in combination with Byte Pair Encoding (BPE) [29], focusing on low-frequency components that represent the overarching shape of the signals. This frequency-domain approach, particularly the strategy of flattening DCT coefficients prioritizing low-frequency information, results in concentrated predictions due to stronger global pattern capture and robust representation against input noise. This approach enhances consistency and stability in action predictions, as evidenced by concentrated class predictions observed in the confusion matrices.
@@
Conversely, OpenVLA employs a discrete binning approach, mapping continuous action spaces to discrete bins directly, thus not capturing global patterns in any sense when evaluated in a zero-shot setting. OpenVLA exhibits sensitivity to variations and input noise, resulting in more evenly distributed probability mass across action classes, as reflected in the broader spread observed in its confusion matrices.
@@ V-1 Differences Between Training Data and Procgen
Dataset Domain Discrepancies: A fundamental obstacle limiting model performance arises from substantial differences between the training datasets and the Procgen evaluation dataset. Procgen environments are procedurally generated, Atari-like 2D games designed to evaluate visual and motor skills of reinforcement learning (RL) agents. Each subdataset within Procgen exhibits diverse environment layouts, tasks, objectives, reward structures, and discrete action spaces, predominantly involving directional movements and specific game-related interactions.
@@
OpenVLA is a vision-language-action model specifically fine-tuned on approximately 970,000 robotics demonstrations from the Open X-Embodiment dataset [22]. These demonstrations are strictly limited to continuous robotic manipulation tasks recorded from third-person camera views.
Pi0 models also leverage robotics trajectory data featuring continuous action spaces. Their training incorporates a subset of OpenX-Embodiment alongside a private dataset containing approximately 903 million timesteps, significantly sourced from single and dual-arm robotic manipulation scenarios [3].
@@
Action Space Awareness: Each model exhibits varying degrees of awareness regarding the action space constraints:
GPT 4x models are explicitly informed through the provided prompts about valid action values and corresponding verbal descriptions.
@@
Pi0 Base and Pi0 Fast rely on training-induced normalization but lack explicit clamping, resulting in a higher chance of invalid predictions outside permissible action ranges.
Proprioceptive State Handling: Procgen datasets lack proprioceptive state information, a critical input for Pi0 models during training. To compensate, zero arrays replace missing proprioceptive states during inference, negatively impacting Pi0 model performance by introducing unnatural, non-representative states. While OpenVLA optionally accepts proprioceptive states, the absence of these in Procgen data inherently limits the richness of information available for inference.
Image Input Variability: Significant discrepancies in image inputs further exacerbate model adaptation difficulties:
@@
View differences: OpenVLA expects a single image input, matching the Procgen input structure, potentially enhancing its adaptability. In contrast, Pi0 models trained on multiple views suffer from performance degradation due to missing additional viewpoints being substituted by zero arrays.
@@ V-2 Difficulty Adapting to Increased Image Complexity
We quantified image complexity through Shannon entropy [30] and Delentropy [25] across 80 episodes over 16 datasets as seen in Figures 37, 22, and 23, revealing distinctive performance patterns. The formula used to calculate the entropies can be seen in Appendix section VII-C:
Figure 22: Entropy measures VS Model Macro Recall.
Figure 23: Correlation matrix for Entropy measures VS Model Macro Recall.
GPT-4 models (4o and 4.1) demonstrated moderate to weak negative correlations (Shannon entropy: -0.409 and -0.183; Delentropy: -0.345 and -0.126, respectively), suggesting poorer performance with increasing complexity.
OpenVLA and Pi0 Base exhibited negligible correlations with entropy metrics (OpenVLA: -0.173 Shannon, -0.063 Delentropy; Pi0 Base: 0.003 Shannon, -0.011 Delentropy), implying relative independence from image complexity.
Pi0 Fast uniquely displayed moderate positive correlations (Shannon and Delentropy: 0.296 and 0.336), indicating improved performance with higher image complexity.
@@ V-3 Impact of VLA Training, Decoding, and Output Processing Techniques
Pi0 Base also targets normalized outputs within [-1,1] without enforced clamping. Its diffusion-based architecture inherently predicts actions “close” to familiar training distributions. This approach yields relatively fewer invalid outputs and moderately effective generalization compared to purely autoregressive models, but remains inferior to OpenVLA.
@@ VI Conclusion
In this study, we systematically evaluated the generalization capabilities of contemporary VLAs and VLMs on procedurally generated discrete-action environments from the Procgen dataset in a zero-shot setting. Our analysis highlighted significant limitations arising from architectural constraints, training paradigms, and input-output biases inherent to the models. The stark domain discrepancy between training data—primarily continuous-action robotics datasets, and general web-scale vision language data—and discrete-action game environments proved to be a critical barrier to effective zero-shot generalization.
We identified notable differences in model behaviors linked directly to their architectures, training strategies, and input/output processing techniques. OpenVLA’s robust action-space clamping technique consistently provided superior generalization, minimizing invalid outputs and exhibiting relative resilience to out-of-distribution scenarios. Conversely, autoregressive models like GPT-4x displayed substantial difficulty in generalizing, especially under complex image conditions, and frequently defaulted to idle or biased action choices. Additionally, Pi0 models showed intermediate performance influenced heavily by their diffusion-based (Pi0 Base) or autoregressive (Pi0 FAST) decoding methods, with Pi0 FAST being notably sensitive to image complexity and unable to restrict the majority of its predictions to a desired output range.
Our findings underscore the necessity for architectural innovations, refined training methodologies, and enhanced output processing techniques to bridge the gap between diverse action domains. Future research should prioritize developing more generalized training datasets that better reflect the variety of potential application environments, alongside methods to adaptively handle different forms of action representations. These advancements hold promise for enabling VLAs and VLMs to operate effectively across an increasingly diverse and unpredictable range of real-world tasks.